OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (3, 4) and has a slope of 2? y = x + Will fan/medal

OpenStudy (anonymous):

@mathstudent55 @paki @iambatman

OpenStudy (ahsome):

First of all, the basic equation of a linear equation is: \[\huge{y=mx+c}\] Where: \[m=Slope\]\[c=y-intercept\]

OpenStudy (ahsome):

The problem is that we don't know what \(c\) is. So we need to find it Another form of expressing a linear equation is: \[\huge{y-y_1=m(x-x_1)}\] Where: \[m=Slope\]\[Point=(x_1,y_2)\]

OpenStudy (anonymous):

\[y-4=m(x-3)?\]

OpenStudy (ahsome):

Exactly. But we know what \(m\) is, right?

OpenStudy (anonymous):

No

OpenStudy (ahsome):

Yes we do. What is the slope?

OpenStudy (anonymous):

Oops 2..

OpenStudy (ahsome):

Its fine :) Now, we know a point, and we know the slope, so lets work from there! \[y-y_1=m(x-x_1)\]\[y-4=2(x-3)\]\[y-4=2x-6\]\[y=2x+2\]

OpenStudy (ahsome):

Sorry, its \(-2\)

OpenStudy (anonymous):

hmm... \[\frac{ y }{ y }=2x+\frac{ 2 }{ y }\]

OpenStudy (ahsome):

Why do you wan't to divide by y?

OpenStudy (anonymous):

Isn't that how it works?

OpenStudy (ahsome):

Nope. You don't need to divide by y :)

OpenStudy (anonymous):

Ugh.

OpenStudy (sepeario):

very very well expressed.

OpenStudy (ahsome):

Thanks, @Sepeario :D